I need help with parent function transformations
f(x)=-4(x+2)3cubed

Answer by checkley77(12844)

f(x)=-4(x+2)3
f(x)=-4(x+2)(x+2)(x+2)
f(x)=-4(x+2)(x^2+4x+4)
f(x)=-4(x^3+4x^2+4x+2x^2+8x+8)
f(x)=-4(x^3+6x^2+12x+8)
f(x)=-4x^3-24x^2-48x-32 ans.